### v2

# A fast, flexible simulation framework for Bayesian adaptive designs - the R package BATSS

Abstract
The use of Bayesian adaptive designs for randomised controlled trials has been hindered by the lack of software readily available to statisticians. We have developed a new software package (Bayesian Adaptive Trials Simulator Software - BATSS) for the statistical software R, which provides a flexible structure for the fast simulation of Bayesian adaptive designs for clinical trials. We illustrate how the BATSS package can be used to define and evaluate the operating characteristics of Bayesian adaptive designs for various different types of primary outcomes (e.g., those that follow a normal, binary, Poisson, or negative binomial distribution) and can incorporate the most common types of adaptations: stopping treatments (or the entire trial) for efficacy or futility, and Bayesian response adaptive randomisation - based on user-defined adaptation rules. Other important features of this highly modular package include: the use of (Integrated Nested) Laplace approximations to compute posterior distributions, parallel processing on a computer or a cluster, customisability, adjustment for covariates and a wide range of available conditional distributions for the response.
Dominique-Laurent Couturier1,†, Rainer Puhr2, Stephane Heritier2,
Thomas Jaki1,3 and Elizabeth G Ryan2
$^{\dagger}:$ Corresponding author,
$^{1}$: MRC Biostatistics Unit, University of Cambridge, UK,
$^{2}$: School of Public Health and Preventive Medicine, Monash University, Australia,
$^{3}$: Faculty of Informatics and Data Science, University of Regensburg, Germany
1 Introduction
Bayesian adaptive designs are increasingly being used to deliver innovative and complex clinical trials with the aim of providing more efficiency and flexibility (e.g., Cellamare et al., 2017; Angus et al., 2020; Denholm et al., 2022; Gajewski et al., 2022). These designs combine features of Bayesian inference that are attractive for trials, such as the ability to incorporate prior information like historical controls and to express treatment effects as probability distributions, with the capability to conduct pre-planned analyses on accumulated data at scheduled time points. This allows for alterations to the study design components, such as sample size, randomisation allocation ratios, and the list of treatments available in the trial, based on the results of these analyses.
Multi-arm, multi-stage (MAMS) designs (Royston et al., 2003; Royston et al., 2011; Magirr et al., 2012; Jaki, 2015) are a type of adaptive clinical trial design that allow for multiple treatments to be compared within the same trial, often to a common control treatment, and aim to select the best treatment(s). This is achieved by performing adaptations such as: early stopping of treatment arms (removal from the study) for efficacy or lack-of-benefit (referred to here as “futility”), early stopping of the trial for efficacy or futility, and updating the randomisation probabilities based on the outcome data, to achieve certain experimental goals (known as “response adaptive randomisation”). Whilst MAMS designs can be conducted in both Bayesian and frequentist frameworks, we focus here on Bayesian MAMS designs which are naturally suited for adaptations as they can be informed by probability statements on treatment effects (which can be updated at each interim analysis), a feature not accommodated by frequentist MAMS designs. Most of the effort in terms of (open source) software development for MAMS designs has been done in the frequentist framework (e.g., nstage (Barthel et al., 2009; Blenkinsop and Choodari-Oskooei, 2019), MAMS (Jaki et al., 2019), rpact (Wassmer and Pahlke, 2023), multiarm (Grayling and Wason, 2020), nstagebin (Bratton, 2014; Choodari-Oskooei et al., 2023), and DESMA (Grayling, 2019)).
Bayesian MAMS and other Bayesian adaptive designs rely on Bayesian inference to adapt trials. At each interim analysis, Bayesian methods combine prior information on model parameters (which might be informed by clinicians, past information, or be vague or non-informative) and the likelihood of parameters given the observed data, resulting in a posterior multivariate probability distribution that summarises the known information about the parameters at a given stage. This distribution allows decision making (such as determining trial or arm successes) to be based on parameter-specific probability statements such as ‘there is a 95% probability that the new treatment is superior to the standard treatment’.
1.1 Monte Carlo Simulations
Whilst traditional designs with a single endpoint may rely on (asymptotic) statistical theory to ensure suitable power and control of type I error, such analytical results are typically not available for complex adaptive designs. Consequently, trial operating characteristics (such as power, type I error, and expected sample size) are typically estimated using Monte Carlo simulations in such cases. Among the many research papers and guidelines describing how to perform and evaluate simulation studies (see, e.g., Morris et al. (2019); White et al. (2023); Ankolekar et al. (2020); Yuxin Ding and Qu (2025)), we focus here on those by
  • the ‘International Council for Harmonisation of Technical Requirements for Pharmaceuticals for Human Use’ (ICH), especially ‘ICH-E20’ the new draft guidelines dedicated to adaptive designs $I_{1}$, (ICH 2025)] and ‘ICH-E9R1’ the draft of the revision of the guidelines dedicated to multiplicity corrections $I_{2}$, (ICH, 2025)],
  • the ‘European Medicines Agency’ (EMA), especially the agency initial reflections on adaptive designs $E_{1}$, (EMA 2007)] and questions and answers dedicated to complex clinical trials $E_{2}$, (EMA, 2022)],
  • the US ‘Food and Drug Administration’ (FDA), especially the agency initial reflections on adaptive Design $F_{1}$, (FDA 2019)] and questions and answers dedicated to complex clinical trials $F_{2}$, (FDA, 2020)],
as they provide detailed guidance on expectations to ensure regulatory acceptance and uptake of complex clinical trials, particularly Bayesian ones. For simplicity, to cite these documents, we will use $X^{p}_{d}$ to refer to page ‘p’ (subscript) of the document ‘d’ (lower script) of source ‘X’, where ‘X’ can either be ‘E’, ‘F’ or ‘I’, respectively corresponding to the EMA, FDA and ICH in the remaining part of this Section. The aforementioned guidelines all mention the need for extensive computer simulations to determine important aspects of the design on which the assessment of design proposals is to be made. Such simulations should be “carefully planned, conducted and reported” ($I_{1}^{5}$) and include, notably,
  • the careful evaluation of the “chance of producing wrong conclusions” ($F^{4}_{1}$) where the probability of erroneous conclusions should be assessed by estimating the frequentist operating characteristics of type I error by computer-intensive simulations ($F^{6}_{1}$, $E^{6}_{1}$). When multiple null hypotheses are being tested, the probability of rejecting at least one of them is often controlled ($F^{19}_{2}$, $E^{5}_{2}$), sometimes referred to as the family-wise error rate (FWER). The EMA considers type I error control to be “a minimum prerequisite for statistical methods to be accepted in the regulatory setting” ($E^{5}_{1}$),
  • a detailed study of other operating characteristics like power and expected, minimum and maximum sample sizes, for example ($F^{11}_{2}$$I^{21}_{1}$),
  • the optimisation, typically by simulation, of the timing and number of interim analyses ($F^{6}_{1}$, $F^{18}_{2}$),
  • a comparison of the above operating characteristics with the ones obtained when considering alternative designs, including conventional (i.e., non-adaptive) ones ($F^{3}_{1}$, $I^{20}_{1}$).
  • sensitivity analyses to different chosen aspects of the design to ensure operating characteristics are robust to deviations of the assumptions made, notably,
    • prior choices ($F^{4}_{1}$, $I^{25}_{1}$), especially when estimating the (frequentist-like) type I error and when considering informative priors via the borrowing of external information (past data, expert opinion). In that case, “it is also important to evaluate the current trial data with no borrowing” ($I^{26}_{1}$).
    • the anticipated timing of interim analyses ($I^{11}_{1}$), related to, for example, changes to Data Monitoring Committee meeting dates,
    • nuisance parameter values ($F^{18}_{2}$, $I^{21}_{1}$), for example,
        the assumed variance for a Gaussian primary endpoint given the predictors (treatment effects and confounders),
        the assumed shape parameter when modelling overdispersed counts with a negative binomial distribution, as well as the average count number in the control group when targeting a given treatment-related fold change,
        the probability of success in the control group when modelling a binary endpoint by means of a binomial distribution and targeting a given treatment-related odds ratio as effect size,
    • drop-out and missingness levels ($I^{20}_{1}$),
    • the use of non-binding futility stopping rules when binding ones were planned, as data monitoring committees sometimes opt for flexibility and might consider the futility stopping criteria as a “guideline that can be deviated from” ($I^{12}_{1}$). Of note, in such cases, the FDA would likely consider the trial “to have failed to provide evidence of efficacy, regardless of the outcome of the final analysis” ($F^{12}_{2}$), hence the importance of such sensitivity analyses.
Other relevant aspects of these guidelines include
  • the suggestion, to mitigate the fact that an infinite number of scenarios are compatible with the selected null and alternative hypotheses ($F^{18}_{2}$, $I^{22}_{1}$), to determine a limited set of scenarios that adequately represent plausible scenarios ($F^{18}_{2}$, $I^{13}_{1}$, $I^{22}_{1}$). This could be achieved, for example, by relying on medical expert opinion ($F^{19}_{2}$). The aim would then be to make sure that the maximum type I error across the grid of scenarios does not exceed the target error rate, for example, 2.5% for a one-sided test ($F^{6}_{1}$), and that the power exceeds the target one in all selected cases.
  • details on the suitable number of Monte Carlo trials per scenario and the way to select seeds per scenario: To achieve a suitable level of precision, it is noted that 100,000 Monte Carlo trials would be sufficient “in most cases” when evaluating the type I error, and that 10,000 simulated trials might be enough to evaluate power ($F^{19}_{2}$, $I^{22}_{1}$). The use of different seeds per scenario is also recommended “to avoid [observing] consistently atypical results across scenarios” ($F^{19}_{2}$).
1.2 Existing software
Grayling and Wheeler (2020) and Meyer et al. (2021) provide recent reviews of software for adaptive designs with the latter focusing on the software potentially available for simulating MAMS designs, and both reviews mention some options for Bayesian designs. These reviews found that simulation of Bayesian adaptive designs generally involved either constructing bespoke code or using standalone simulator software (such as FACTSTM (FACTS Development Team, 2023), HECT (Thorlund et al., 2019), or EAST® (Cytel Inc., 2020)).
The advantage of these standalone simulators is that they do not require the user to perform any programming and usually have been optimised to perform the trial simulations quickly. However, they are often limited in the designs that can be explored (types of endpoints and models used, adaptations permitted, stopping rules implemented), may only provide pre-defined operating characteristics, may not allow the user to alter the simulation control parameters (i.e., length of the warm-up/burn-in phase, number of samples, number of chains (for the sampling algorithm)) or provide convergence diagnostics for sampling algorithms, may require a commercial licence, and/or have graphical user interfaces that hinder reproducibility.
Script-based software/packages offer several advantages over stand-alone simulators, the main advantages being that they permit reproducibility, and are generally open-source. However, there is a lack of script-based packages available for simulating Bayesian adaptive designs. There are a small number of R packages for simulating two-arm, Bayesian group sequential designs (gsbDesign (Gerber and Gsponer, 2016), goldilocks (Hickey et al., 2022)) which allow early stopping for efficacy or futility, and the adaptr (Granholm et al., 2022) package can be used to simulate Bayesian MAMS designs. In general, these packages can only be used for certain outcomes, are limited in the types of designs or adaptations that can be implemented (often with little flexibility in the choice of adaptive rules that can be used) and may only offer simple models for the outcome.
For those who do not have access to closed-source simulator software and/or find existing packages unsuitable for their trial design needs, simulations will need to be performed using custom-written code. Whilst bespoke code is tailored to meet the needs of a project, it may only be used for a specific trial design, and may lack the generalisability to be re-used for future projects (Meyer et al., 2023). Such code may also lack validation. Additionally, programming project-specific code may also be an inefficient use of statisticians’ time and lead to delays in starting a clinical trial.
Therefore, the development of open-source trial simulation software that is reliable, offers good flexibility in the designs and adaptations available, can cope with different types of endpoints (binary, continuous, count, time-to-event) and is fast is highly desirable.
1.3 Aim
The aim of this work is to provide an open-source, script-based, flexible structure (via an R software package) that can be used as a general design tool for the fast simulations of Bayesian adaptive trials and that would enable the efficient evaluation of design characteristics across a wide range of scenarios and number of simulations, as required by the EMA, FDA, and ICH guidelines mentioned in Section 1.1.
We will illustrate how our Bayesian Adaptive Trials Simulator Software (BATSS) R package can be used to define and evaluate the operating characteristics of a Bayesian adaptive trial for different types of endpoints given some common adaptations: stopping arms (or the entire trial) for efficacy or futility and response adaptive randomisation. To permit generalisability, we use a modular approach where the user can select the adaptations that they wish to include and specify their own adaptation rules through user-defined R functions. We provide several examples of such functions to assist the user.
This package is targeted at statisticians with some experience in adaptive designs, knowledge of Bayesian statistics, who are familiar with R, and who might not have access to commercial clinical trial simulator software or may not want to write their own custom-code to simulate a design.
This paper details the implementation, methods and functions of the BATSS R package. In Section 2 we provide a brief overview of the methods used in the BATSS package for approximating the posterior distribution. Section 3 provides an overview of the BATSS framework, functions and workflow. Two case studies, respectively considering a negative-binomial and binomial endpoint, are introduced in Sections 3 and 4, along with an illustration of how these designs would be implemented using the BATSS package and their operating characteristics. We conclude with a discussion on the use of the BATSS package and directions for future extensions of this package.
2 Laplace approximations for the posterior distribution
In the Bayesian generalised linear models (GLM) considered in this work, the response vector of $n$ observations, $\mathbf{y}=[y_{1},...,y_{n}]^{T}$, is assumed to follow a density function depending on parameters $\bm{\mu}$} and $\phi$
$\mathbf{y}\sim f(\mathbf{y}\ |\ {\bm{\mu}},\ \phi)$
where $f(\cdot)$ is a density distribution belonging to the exponential family (or similar), $\bm{\mu}$} denotes the expectation of $\mathbf{y}$, and $\phi$ is an optional nuisance parameter. The exponential family encompasses a wide range of distributions like the binomial distribution for binary endpoints, the Poisson and negative binomial distribution for count data, and the beta, Gaussian and gamma distributions for continuous endpoints. In GLMs, the response mean is linked to the model predictors through a link function:
${\bm{\mu}}=g(\mathbf{X}{\bm{\beta}})$
where $\mathbf{X}$ denotes the ($n\times p$) predictor design matrix, and $\bm{\beta}$} denotes the regression parameter vector whose interpretation depends on the chosen link function. The expit function ${\bm{\mu}}=\exp(-\mathbf{X}{\bm{\beta}})^{-1}$ is often used for endpoints bounded between 0 and 1 (like binary or beta endpoints), leading to $\bm{\beta}$} coefficients corresponding to odd ratios. The exponential function ${\bm{\mu}}=\exp(\mathbf{X}{\bm{\beta}})$ is typically used for strictly positive endpoints (like counts), leading to $\bm{\beta}$} coefficients corresponding to fold changes. The linear model is a special case of GLM in which $f(\cdot)$ corresponds to the Gaussian distribution and ${\bm{\mu}}=\mathbf{X}{\bm{\beta}}$ (identity link function).
In the frequentist framework, the GLM parameter vector combining the $\bm{\beta}$} regression parameters and optional nuisance parameter $\phi$, ${\bm{\theta}}=[{\bm{\beta}},\phi]^{T}$, is estimated by maximising the likelihood function of the parameters given the data. In the Bayesian framework, the parameter vector is assumed to follow a (multivariate) prior distribution ${\bm{\theta}}\sim\pi({\bm{\theta}})$. Bayesian inference then asserts that the posterior probability distribution of the fitted model parameters is proportional to the likelihood function multiplied by the prior distribution of the parameters.
$\pi({\bm{\theta}}\ |\ \mathbf{y},\mathbf{X})\propto f(\mathbf{y}\ |\ {\bm{\theta}},\mathbf{X})\ \pi({\bm{\theta}})$
For each iteration of the Monte Carlo simulations (e.g., 1…M=10,000 trials) that are performed to understand the operating characteristics of a Bayesian adaptive trial design, the posterior distribution must be computed (often on multiple occasions within a simulated trial for each interim analysis that is performed). Thus, fast computations of the posterior distribution are essential. Calculation of the posterior distribution can often involve high-dimensional integration for which closed-form expressions do not exist and so numerical techniques are needed.
Sampling-based methods, such as Markov Chain Monte Carlo (MCMC) methods, have commonly been used for estimating the posterior distribution and related quantities for Bayesian inference. These methods require large numbers of MCMC samples which can be computationally demanding (and sometimes prohibitive), and necessitate both the assessment of convergence of the underlying Markov chain to stationarity as well as the assessment of convergence of Monte Carlo estimators to population quantities (Roy, 2020; Robert and Casella, 2010). To circumvent these problems, we use a different approach, which is deterministic and potentially much faster.
2.1 Integrated Nested Laplace Approximations (INLA)
Laplace approximation provides an analytical method to estimate the posterior distribution of parameters in complex models by assuming that the posterior distributions can be approximated by Gaussian densities centered at their mode (or maximum). The calculation of integrals related to the posteriors are therefore greatly simplified, typically leading to computational gains, especially for complex models. R has several implementations of Laplace approximations for Bayesian inference, such as in the TMB (Template Model Builder), LaplacesDemon and mgcv packages, for example.
BATSS relies on the Integrated Nested Laplace Approximations (INLA) framework (Van Niekerk et al., 2023; Brown et al., 2021; Rue et al., 2009) that provides approximate analytical posteriors by assuming that the model follows a specific structure – the latent Gaussian model – and uses sparse matrices (Rue et al., 2009). Latent Gaussian models represent a wide class of commonly-used statistical models, including generalised linear (mixed) models, spatio-temporal models, survival and joint longitudinal-survival models. Details about the modelling structure and default priors can be found in (Couturier et al., 2025).
INLA is implemented in the R software using the R-INLA package (R-INLA Project, 2023). INLA has proved to be popular in many applied areas, including spatial statistics, ecology, and epidemiology (e.g., Alene et al., 2022; Pimont et al., 2021) and can potentially be used to perform inference for many types of clinical trial primary outcomes.
In addition to the flexibility provided by the broad range of models that INLA can accommodate, the computational efficiencies that it offers for performing Bayesian inference were a strong motivating factor for incorporating INLA into our Bayesian adaptive trial simulator program. In many situations, INLA methods can provide faster and as accurate inference as MCMC methods (Rue et al., 2017). Indeed, Yimer and Shkedy (2021) noted, “The computational time for INLA is in seconds whereas the computational time for JAGS and Stan is in minutes.” Parallel processing is also a built-in feature of INLA that further enables computationally efficient Bayesian inference, independent of the specific model that is implemented (Antónia et al., 2019). In the context of adaptive designs, evidence regarding the computational efficiency of INLA has recently emerged. Hosseini et al. (2024) used INLA to ensure the feasibility of the extensive simulations required to design a Bayesian adaptive trial to evaluate novel mechanical ventilation strategies in patients with acute respiratory failure. Speed was particularly an issue for this study based on a hierarchical proportional-odds model. The authors explained that the simulation program would have taken years using R Stan (Stan Development Team, 2023) and MCMC, instead of 20 days with INLA.
3 The BATSS R package
The aim of the BATSS R package (Couturier et al., 2025) is to provide a user-friendly tool for the fast simulation of Bayesian adaptive trial designs so that a design’s operating characteristics can be determined. It is intended that the BATSS package can be used for phase II-IV trials with interim analyses for stopping arms or the trial and alterations to the sample size or allocation ratio. As a first step, we focus our attention on MAMS designs with a common control arm (where we are primarily interested in pairwise comparisons to the control), implemented in the Bayesian framework.
BATSS is highly customisable: user-defined functions are used to simulate the outcome variable, treatment arm allocations, additional covariates, and also for the adaptation rules. We use a modular structure so that users can select the desired types of adaptations and level of complexity of the design. Installation instructions for BATSS can be found in Appendix A.
In this work we define “intervention” arms as the experimental treatments that are under investigation (i.e., the non-control arms); “treatment” arms refer to all of the arms included in the study, both control and intervention. We define “active” treatment arms as a treatment arm (intervention or control) that is still in the trial and has not been dropped. We also use the terms “adaptive analyses” and “interim analyses” interchangeably. “Looks” of the data include interim/adaptive analyses and a final analysis at the maximum sample size (if the trial does not stop early). The “target parameters” are the model parameter estimates that correspond to the treatment effects of interest.
3.1 Case Study 1 - overdispersed count outcome
Here we consider a case study to illustrate the different components of the BATSS package. This example is inspired by the CHANGE-MS study (Curtin et al., 2016), which was a multi-arm, phase II randomised controlled trial that investigated the use of temelimab for the treatment of relapsing-remitting multiple sclerosis. This example trial consists of four treatment arms: 3 interventions - low, medium and high doses of temelimab (referred to as arms “A”, “B” and “C”, respectively) and a control. The outcome is an overdispersed count outcome (cumulative number of new active lesions) that is modelled by a negative binomial distribution (Curtin et al., 2016; Hartung et al., 2022), where a lower value indicates a more favourable response.
[…]
3.2 Overview of the BATSS framework
The overarching function that is used to simulate the trial design is batss.glm, which can be used for designs where the data is modelled by a generalised linear model (GLM). Figure 1 provides a depiction of the inputs and output generic functions of batss.glm. This function takes a number of arguments that the user can easily alter, and several defaults are provided; these are described in detail Table A1 in Appendix A and in the help function for batss.glm.
Figure 1: Summary of the inputs and outputs of the batss.glm function, demonstrating its modular structure.
3.3 batss.glm workflow
Algorithm 1’s pseudocode describes the process that the batss.glm function performs when simulating a Bayesian adaptive trial design. A workflow description is also provided in Appendix A.
It is important to note that, if the trial does not stop early, the final analysis is performed at the maximum sample size, using the user-defined efficacy and futility rules for each arm. The maximum sample size is specified for the entire trial (rather than for each arm) and so remaining treatment arms may be enriched with larger sample sizes if some arms are terminated early.
If RAR is selected for the design, it is performed at each adaptive analysis point (after the burn-in period of $m_{0}$ participants). The user can select whether to implement efficacy or futility stopping from the first adaptive analysis onwards, or to start efficacy or futility stopping at a later analysis timepoint (through the delta arguments of the batss.glm function).
Algorithm 1 A simplified version of the batss.glm simulation workflow. Relevant arguments of the function appear in blue.
1:for $i=1,...,{\color[rgb]{0,0,1}\texttt{R}}$ do
2:  ## Initialisation
3:  Set $m=m_{0}$ # 1st set of patients [interim=list(recruited=c(m0, ...))]
4:  Set $prob=\color[rgb]{0,0,1}[\texttt{prob0}]$ # initial randomisation probabilities
5:  Set $n_{curr}=m$ # current sample size
6:  while Any intervention arms are active and $n_{curr}\leq{\color[rgb]{0,0,1}\texttt{N}}$ do
7:   Simulate $m$ sets of predictors [var, var.control]
8:   Simulate $m$ responses [model, var, var.control, link]
9:   Fit model to the data using INLA [model, family, link]
[…]
11:   for $k$ in no. of active intervention arms do
12:    Assess efficacy stopping rules [eff.arm, eff.arm.control]
13:    Assess futility stopping rules [fut.arm, fut.arm.control]
14:   end for
15:   Assess trial stopping rules [eff.trial, fut.trial]
16:   if All intervention arms inactive or trial stopping rules have been met then
17:    BREAK
18:   end if
19:   if RAR $\neq$ NULL then
20:    define randomisation probabilities [RAR, RAR.control, delta.RAR]
21:   end if
22:   Update $m$, $n_{curr}$, and probprob
23:  end while
24:end for
3.4 The batss.glm function - syntax
Case Study 1 may be simulated using the following syntax:
sim = batss.glm(
model = y~treatment,
var = list(y = rnbinom,
treatment = alloc.balanced),
var.control = list(y = list(size = 1/2)), # shape
family = "nbinomial",
link = "log",
beta = log(c(4, 1, 1, 0.4)),
which = c(2:4),
alternative = c("less"),
R = 10000,
N = 260,
interim = list(recruited=c(100,140,180,220)),
prob0 = c(control = 1, A = 1, B = 1, C = 1),
delta.eff = 0,
eff.arm = eff.arm.infofract,
eff.arm.control = list(b.eff = 0.009, p.eff = 3),
delta.fut = log(0.8),
fut.arm = fut.arm.simple,
fut.arm.control = list(b.fut = 0.2025),
RAR = NULL,
RAR.control = NULL,
H0 = FALSE,
computation = "parallel",
mc.cores = detectCores()-1,
extended = 1)
Here we use some pre-defined functions for the treatment allocations (alloc.balanced), efficacy (eff.arm.infofract) and futility (fut.arm.simple) stopping.
The operating characteristics for the above simulated scenario can be obtained using:
summary(sim, extended = TRUE)
Figure 2 shows the probability of stopping the trial at each look under the assumed parameter values as well as the expected trial sample size. Figure 3 displays the sample size of each simulated trial per group with, for the target parameters, points colour-coded and shaped according to the decision and timing when the trial stopped. Figure 4 shows the parameter estimates (posterior mean) against the sample size for each arm and the decision at the time of stopping the arm for the simulated design. These plots can be obtained using:
plot(sim, which = 2:4)
where which=1 produces the boxplot of the total and per group sample sizes of the simulated trials.
Figure 2: Barplot of the probability of stopping (y-axis) at each look (x-axis) as observed in the simulated samples. The vertical blue dashed line shows the estimate of the expected sample size (ESS).
Figure 3: Violin plot (and boxplot) of the simulated sample sizes per treatment arm for each trial and corresponding expected sample sizes. For target parameters, these points are colour-coded and shaped by the decision and timing of when the arm stopped.
Figure 4: Scatterplot of target parameter estimates (posterior mean) versus sample size for each intervention arm (separate plots) at the time of stopping the arm, along with the decision at the time of stopping for the simulated scenario for Case Study 1 across the 10,000 simulated trials. The timing of the stop is also provided (“Early stopping” = stopping at an interim analysis).The horizontal blue lines show the assumed target parameter values.
A fuller description of Case Study 1, along with the full code to simulate the example, output and interpretation of the output can be found in Appendix B.
We now briefly describe the arguments of the batss.glm function (and provide a more detailed description in Table A1 in Appendix A). A vignette is also available in Couturier et al. (2025).
3.5 Defining the model - types of outcomes, data simulation and model fitting
R-INLA provides inference for many types of GLMs and allows for several different families for each data type, such as:
  • Binary data: binomial, beta-binomial
  • Count data: Poisson, zero-inflated Poisson, negative binomial
  • Continuous data: Gaussian, Weibull, exponential, lognormal, gamma
  • Bounded data: beta
This offers additional flexibility over many existing adaptive design simulation packages, which are often limited in the distributions that they support.
Our framework can handle general cases in which the response is a function of the treatment arm allocation and additional predictors, with or without interaction terms, and in which the target parameters can include both the main and/or interaction parameters. However, for illustrative purposes, we consider simpler models in this paper and work under the assumption that the user-specified model is a GLM with a linear predictor ($g(\mu)$) of the form
$g(\mu)=\mathbf{X}\bm{\beta}=\beta_{0}+\sum_{k=1}^{K}\,\beta_{k}\,I_{k}$ (1)
where $\mathbf{X}$ denotes the model design matrix and $\bm{\beta}=[\beta_{0},...,\beta_{K}]^{T}$ is the corresponding parameter vector, and $I_{k}$ is an indicator for treatment arm $k$ (of which there are $K$ interventions to compare to control; $K+1$ arms in total included in the study).
In the batss.glm function, the user specifies the family type (family = ) and link function (link = ) that are to be used when fitting the model (these must correspond to the options provided in R-INLA). The user must also specify a model object (model = ) indicating both how to simulate and fit the response (i.e., the same model is used for both purposes). Predictor variables (such as the treatment allocation (factor)) are simulated according to user-specified functions defined in the var and var.control arguments and the user-specified model is used to generate the linear predictor value ($\mathbf{X}\bm{\beta}$). This is then used to simulate the response, using the inverse link function and the user-specified outcome distribution and parameters (defined in the var and var.control arguments).
Additional predictors can be added to the model by including their variable name in the model formula, as well as their data generating process and related parameter values in the var and var.control arguments. Note that it is possible to generate dependent predictors by specifying a multivariate data generating process in the var argument.
For Case Study 1, a dispersion (shape) parameter with a value of 0.5 was assumed using the following parameterisation for the variance:
$Var(Y)=\mu+\mu^{2}/\phi$ (2)
(i.e., $\phi=1/2$). This is a nuisance parameter that has to specified and this is done via the command var.control= list(y = list(size = 1/2)).
The user can specify the assumed true parameter values (corresponding to those in equation 1) via the beta argument (on the link function/linear predictor scale). The interpretation of the elements of this vector matches the one we would have in a fit obtained via the lm or glm R functions, and therefore depends on the way predictors are generated and contrasts of factors are defined. For Case Study 1, using the model specified in (1), assuming treatment or dummy contrasts (Logan, 2010) are used for the factor ‘treatment’ with ‘control’ as the reference group and the use of the log link, then the first value of beta corresponds to the mean response in the control treatment (the intercept) on the log scale, and the second, third and fourth elements correspond to the logarithm of the relative risks for treatments A, B and C compared to control. If we assume a mean count of 4, 4, 4 and 1.6 for the control and groups A, B and C, respectively (on the original scale), we would then have the model $g(\mu)=log[E(Y)]=log(4)+log(1)\,I_{A}+log(1)\,I_{B}+log(0.4)\,I_{C}$ (where $I_{A}$, $I_{B}$ and $I_{C}$ are indicator variables for treatment arms A, B and C, respectively) and $\bm{\beta}=[log(4),log(1),log(1),log(0.4)]^{T}$. The user then specifies which of those parameters correspond to the parameters of interest (known as the “target” parameters), via the which argument. In our model, we are interested in the incremental effects of treatments A, B and C with respect to the control so that the target parameters are the second, third and fourth elements of beta and which=c(2:4).
3.6 Design specification - interim analyses and adaptations
The direction of the alternative hypothesis is specified via the alternative argument for each target parameter (one-sided hypothesis testing), where alternative="less" in this instance as a lower value indicates a more favourable response. The maximum sample size of 260 is provided via the N argument.
The user provides the timing of the interim analyses in the interim argument. Currently, batss.glm assumes that the outcome data is immediately available and that interim analyses can only be performed based on the number recruited (interim = list(recruited = )). Thus, here, interim analyses based on the number recruited is equivalent to interim analyses based on the number with complete follow-up, i.e., the same amount of information will be available under both monitoring schedules. The interim analyses can be equally-spaced or unequally spaced.
The batss.glm function supports the following types of adaptations: stopping arms for efficacy (also known as “graduation” of arms), stopping (dropping) arms for futility, stopping the entire trial early to declare one or more interventions superior to control, stopping the entire trial early due to statistical futility (e.g., lack-of-benefit), sample size re-estimation, and response-adaptive randomisation (RAR). Fixed allocation ratios can also be used instead of RAR (RAR = NULL).
As a design’s operating characteristics are strongly influenced by the chosen efficacy and futility stopping rules as well as the chosen method to set the group allocation probabilities, flexibility in defining these adaptations is of crucial importance. BATSS therefore allows these rules to be user-defined, thus allowing great flexibility.
The efficacy and futility stopping rules (for the intervention arms) are provided via user-defined functions in the eff.arm and fut.arm arguments, respectively (with input parameters provided in eff.arm.control and fut.arm.control). Efficacy and futility stopping rules of the trial at a given interim analysis are respectively indicated via the eff.trial and fut.trial arguments that take the target parameters results as input. This, for example, allows the trial to stop once one intervention arm has been found efficacious. Several example functions are provided in the BATSS vignette (Couturier et al., 2025). We defer their detailed explanation to the case studies (Section 4, and Appendices B and C).
Initial randomisation probabilities/allocation ratios, prior to RAR implementation (i.e., during the “burn-in” period before the first interim analysis), are provided in prob0. These probabilities are used throughout the trial if RAR is not implemented. Here we chose equal randomisation throughout the trial, therefore prob0 = c(control = 1, A = 1, B = 1, C = 1) (note that one could also set prob0 = c(control = 1/4, A = 1/4, B = 1/4, C = 1/4)). These probabilities will be re-scaled if a treatment is dropped.
The user-defined RAR rules are provided via the RAR argument (with input parameters for the RAR function given in RAR.control). The user-defined functions permit restrictions of RAR, such as allocation rules with fixed proportions to control or minimum or maximum allocation probabilities to be implemented. This will be illustrated Case Study 2 and several examples are provided in the BATSS vignette.
These different adaptations typically rely on the posterior probability of being greater or smaller than the clinically important difference indicated in the delta arguments – respectively delta.eff and delta.fut for the efficacy and futility arm stopping rules, and delta.RAR for the RAR probabilities – which is estimated at each look for each target parameter.
We provide some examples of functions for the adaptations in Section 4 and Appendices B and C, and describe the required inputs for these functions in Table 1 and Table A2 in Appendix A.
3.7 BATSS ingredients
The batss.glm function calculates a number of “ingredients” or derived quantities at each interim analysis which can be used as inputs to the user-defined functions for the adaptations. These are summarised in Table 1 (full description given in Table A2 in Appendix A, and in the vignette (Couturier et al., 2025)).
Quantity Description
active A logical vector indicating which treatment arms are still active/remaining in the trial.
posterior Posterior probability of efficacy (futility) of an “active” intervention, incorporating the deltas (clinically important effects).
n A vector providing the number of participants that have been recruited in each treatment arm; includes control and intervention arms (both active and dropped arms).
ref A logical vector indicating which treatment arm is the reference (control).
prob Randomisation probabilities (for the active arms). The output from the RAR function (unnormalised). If RAR=NULL in batss.glm then prob=prob0.
m number of subjects recruited since the last interim analysis
Table 1: Description of the common quantities that batss.glm calculates at each interim analysis that can be used as inputs to the user-defined functions for adaptations
3.8 Parallel and cluster computing
In BATSS, the number of simulated trials is provided by means of the argument R, either with a scalar denoting the number of trials (in that case seeds 1 to R will be used, e.g., 1 to 10,000 for our example above) or with a vector of seeds to use to generate each trial’s data (in that case the number of trials will equal the length of that vector).
The BATSS package takes advantage of computing tools for embarrassingly parallel workloads which allows the user to split the burden of the independent Monte Carlo trials between multiple cores on single computers or multiple cores/Central Processing Units (CPUs) of different computers/clusters:
  • When performing computation on the same computer, the user can use parallelisation tools of the parallel package by setting the argument computation to ‘parallel’ and by specifying the number of child processes – Monte Carlo trials in this context – to be run in parallel via the argument mc.cores.
  • To split the computational burden between different computers and/or the CPUs of one or several clusters, the user can split the planned list of seeds into non-overlapping sets. Each set will then be assigned to an available resource (like a CPU on a cluster), either by hand or more efficiently via a script (like an sbatch on clusters using Simple Linux Utility for Resource Management (SLURM) as the job manager), by a call of the batss.glm function using the set of seeds of interest through the R argument and saving the results of each trial (setting argument extended >> 0). The batss.combine function, taking as input the location of the R data files containing the outputs corresponding to each set of seeds, will then be used to aggregate results. (Note that batss.combine checks if the sets of seed overlap or if the calls to batss.glm were different).
We provide some exemplar code in a vignette demonstrating how to run the batss.glm function on a cluster and how to use the batss.combine function (see Couturier et al. (2025)) .
3.9 Output
The output generated by batss.glm is an S3 object of class batss (with type = glm), with available print, summary and plot ‘methods’. A number of different operating characteristics are provided in the output generated by batss.glm.
The summary method (obtained by using the summary function on the output of batss.glm) first prints a summary of the design and inputs to the batss.glm function. Then, in its “short” variant (obtained by opting for extended = FALSE), it indicates the probability of declaring efficacy or futility for each intervention arm, for at least one intervention arm, as well as for all intervention arms for each hypothesis of interest (i.e., under the null and/or alternative hypotheses if argument H0=TRUE, for example). Note that, as a diagnostic tool, the probability of simultaneously meeting both the efficacy and futility criteria is also provided if this highly undesirable situation occurred for any arm in any simulated trial. In its “long” variant (obtained using the summary function with extended = TRUE option), the method additionally provides: a breakdown of when the efficacy or futility stopping occurred (early = at an interim analysis; last = at the maximum sample size); summaries of the sample sizes for each arm and overall (mean, standard deviation, median and 10th and 90th percentiles); and a summary of the different scenarios that occurred in terms of the decisions for each intervention arm (no decision, efficacy decision, futility decision, both efficacy and futility criteria met (this is only displayed for simulated designs where this undesirable behaviour has occurred)) and how frequently those different combinations occurred across the Monte Carlo trials (and what proportion of those had early decisions made where the trial stopped at an interim analysis).
By using the extended argument in the batss.glm function, users can specify what they want to be stored in the output: 0 only saves the overall summary operating characteristics, 1 adds the trial results with a breakdown by look/analysis point to them, and 2 includes all previous results plus the R simulated datasets.
Plots of the posterior mean of the target parameter (i.e., treatment effect) estimates versus the sample size at the last look performed for that Monte Carlo trial, along with the decision (efficacy, futility, none, both), for each intervention arm, can be obtained using the plot function with the which = 4 option (demonstrated in Figure 4). Plots of the distributions of the sample sizes per group (across the Monte Carlo trials) can also be obtained using the which = 3 option in the plot function (demonstrated in Figure 3), as well as using the which = 1 option.
We now present a second case study to further illustrate the use of the batss.glm function and how to summarise its output.
4 Case Study 2 - binary outcome
Here we demonstrate how to use the batss.glm function to simulate a Bayesian MAMS trial design with a binary outcome. This example is motivated by the Personalised Immunotherapy Platform (PIP) trial (Lo et al., 2022), which is a phase II randomised controlled trial that is currently under development that will compare immunotherapy combinations for advanced melanoma patients.
[…]
Interim analyses will be performed once 60 participants have completed follow-up, and then every 12 participants thereafter, up to a maximum sample size of $N=216$ (leading to a total of 14 analyses). A burn-in period of 60 participants will be used during which equal randomisation will be implemented (1:1:1:1:1:1). After this, Bayesian RAR will begin. The randomisation probabilities are calculated using a variation of Trippa et al. (2012) and Cellamare et al. (2017) where the allocation probabilities at look $j$ are proportional to:
  • control (reference) arm:
    $\frac{\exp(\text{max}(n_{1,j},n_{2,j},...,n_{5,j})-n_{0,j})^{\nu}}{K_{j}}$ (3)
    with $(n_{1,j},n_{2,j},...,n_{5,j})$ being a vector of the number of participants recruited to each intervention arm (arms B-F) at look $j$, $n_{0,j}$ is the number of participants recruited to the control arm at look $j$, $K_{j}$ = number of active intervention arms ($K_{j}+1$ is the number of active arms including control) at look $j$ (i.e., number of arms still in the study at that point in time), $\nu$ is a tuning parameter.
  • intervention arms (where $k>0$) :
    $\frac{P(\beta_{k}>\delta_{r}|\ \mathbf{y},\mathbf{X})^{h}}{\sum^{K_{j}}_{l=1}P(\beta_{l}>\delta_{r}|\ \mathbf{y},\mathbf{X})^{h}}$ (4)
    with $h=\gamma(\sum^{K}_{k=0}n_{k,j}/N)^{\eta}$, where $n_{k,j}$ is the number of participants enrolled in arm $k$ at look $j$ ($K$ is the total number of intervention arms initially included in the study, not including the control; $k=0$ is control), and where $\mathbf{y}$, $\mathbf{X}$ respectively correspond to the vector of responses and model design matrix. Here $\beta_{k}$ is the treatment effect for intervention arm $k$ (log-odds scale) and $\gamma$ and $\eta$ are tuning parameters (see Gotmaker et al. (2019)). $\delta_{r}$ is the (RAR-related) clinically meaningful treatment effect value, and for this example we set $\delta_{r}=0$.
These RAR functions (equations 3 and 4) can be written in R as:
prob.trippa = function(posterior, n, N, ref, active, gamma,eta,nu){
g = sum(active) # (K_j+1 in the above)
h = gamma*(sum(n)/N)^eta
# vector of allocation probabilities:
prob = rep(NA,g)
# reference/control arm allocation - Equation 3:
prob[1] = (exp(max(n[!ref])-n[ref])^nu)/(g-1)
# targets/interventions (that haven’t been dropped) - Equation 4:
prob[2:g] = (posterior^h)/(sum(posterior^h))
prob
where the arguments in orange are BATSS ingredients (Table 1), and those in red are items from the respective user-provided ‘.control’ arguments in the batss.glm function. This same convention applies throughout this section.
We would then input these probabilities into a treatment allocation function, such as:
# specify group allocation function
group.fun = function(m, prob){
prob = abs(prob)/sum(abs(prob)) # (normalise probabilities)
factor(sample(names(prob), m, replace=TRUE,
prob = prob), levels=names(prob))
Using these RAR functions (equations 3 and 4), the allocations to the control arm are protected so that sufficient power can be obtained for the comparisons of interest.
Stopping arms for efficacy is not permitted at the interim analyses in this case study as the main objective of the adaptations is to drop poorly performing arms.
[…]
$P(\beta_{k}>\delta_{f}|\ \mathbf{y},\mathbf{X})\ \ <\ \ b_{f}$ (5)
where $\beta_{k}$ is the treatment effect for intervention $k$ (on the log-odds scale), and $\delta_{f}$ is the minimum clinically important treatment effect (on the log-odds scale), i.e.,
$\delta_{f}=log\left(\frac{(\pi_{0}+0.1)/(1-(\pi_{0}+0.1))}{\pi_{0}/(1-\pi_{0})}\right)$
($\pi_{0}$ is the control arm ORR and $\pi_{0}+0.1$ is the minimum desired ORR in an intervention arm).
We begin by using a fixed cut-off of the futility boundary of 0.1, i.e., $b_{f}=0.1$ (Case Study 2.1). This futility stopping rule (equation 5) can be written in R as:
futility1.arm.fun = function(posterior, b.fut){
posterior < b.fut
We will also investigate a stopping rule where the futility boundary increases as a function of the existing number of participants that have been randomised (Case Study 2.2):
$P(\beta_{k}>\delta_{f}|\ \mathbf{y},\mathbf{X})\ \ <\ \ b_{f}\left(\frac{\sum^{K}_{k=0}n_{k,j}}{N}\right)^{p_{f}},$ (6)
where $b_{f}$ is the futility (lack-of-benefit) boundary at the maximum sample size, $\sum^{K}_{k=0}n_{k,j}$ is the total number of participants enrolled at look $j$, and $p_{f}$ is a tuning parameter that dictates the shape of the function (Gotmaker et al., 2019). That is, equation 6 is an increasing power function of the fraction of information available in the trial at that stage. Here we use an increasing cutoff on the posterior probability of observing a meaningful treatment effect as we should have more evidence as the trial progresses; an arm stops for futility if this boundary is crossed. This futility stopping rule (equation 6) can be written in R as:
futility2.arm.fun = function(posterior, n, N, b.fut, p.fut){
posterior < b.fut*(sum(n)/N)^p.fut
An intervention may be declared effective at the end of the trial (at the maximum sample size) if the posterior probability of efficacy is above a certain (high) threshold:
$P(\beta_{k}>\delta_{e}|\ \mathbf{y},\mathbf{X})\ \ >\ \ 1-b_{e}$ (7)
Here we use a value of $\delta_{e}=0$ for the final look (here $\delta_{e}$ would also represent a clinically important treatment effect on the log-odds scale; this is the default value in batss.glm()). The value of $b_{e}$ is typically small and chosen to allow good control of type I error; we aim to control the pairwise one-sided type I error at $5\%$. The efficacy rule (equation 7) can be written in R as:
efficacy.arm.fun = function(posterior, b.eff){
posterior > 1-b.eff
# where delta.eff needs to have NAs to not evaluate efficacy at interims
The trial will only stop early if all intervention arms have been dropped for futility; otherwise the trial will proceed to the maximum sample size of $N=216$ and evaluate efficacy and lack-of-benefit for the remaining arms at the end of the trial. Note that care should be taken when constructing the efficacy and futility stopping rules as it is possible to simultaneously meet the efficacy and futility criteria in certain situations (see Appendix C, Case Study 2.2).
[…]
The pairwise one-sided type I error was calculated as the proportion of simulations that falsely declared efficacy when the intervention was assumed to have the same ORR as the control (i.e., log OR = 0). The (pairwise) power was calculated as the proportion of simulations that declared efficacy for an intervention when it was assumed to have a large effect (i.e., $20\%$ absolute improvement in the ORR compared to the control; as in the “alternative” scenario).
6-month ORR
Scenario Arm A (control) Arm B Arm C Arm D Arm E Arm F
Global Null 0.4 0.4 0.4 0.4 0.4 0.4
Alternative 0.4 0.4 0.4 0.5 0.7 0.7
Table 2: Scenarios explored for PIP trial case study (Case Study 2)
Here we have focused on providing the user-defined functions for performing the adaptations. The full description of the batss.glm() call for simulating the “Alternative” scenario for Case Studies 2.1 and 2.2 can be found in Appendix C, following the logic described above.
4.1 Results
The output from the summary command can be found in Appendix C, along with an explanation of the output.
4.1.1 Case Study 2.1
Table 3 shows the proportion of simulations that declared each intervention arm to be effective (at the maximum sample size/end of the trial), the proportion of simulations that each intervention arm was stopped early to declare futility, and the average sample size per arm and overall sample size, for each scenario for Case Study 2.1. Figure 5 shows the probability of stopping at a given look and confirms that early stopping – only occurring when all arms are found futile – is rare. The distribution of sample sizes for each treatment arm can also be examined in Figure 6 for the “Alternative” scenario. A summary of the target parameter estimates against the sample size achieved for a simulated trial is provided in Figure 7.
[…]
Conclusion Sample size
Group Efficacy(1) Futility(2) Average(3) Stand. Dev.(4)
Under the global null hypothesis:
Control (Arm A) - - 49.13 17.31
Arm B 0.044 0.604 28.36 17.97
Arm C 0.044 0.609 27.91 17.81
Arm D 0.046 0.611 27.62 17.68
Arm E 0.046 0.608 28.13 17.93
Arm F 0.045 0.605 28.09 17.98
At least 1 0.156 0.890 - -
All 0.001 0.313 189.24 48.61
Under the alternative hypothesis:
Control (Arm A) - - 49.89 9.46
Arm B 0.046 0.516 20.59 9.76
Arm C 0.040 0.518 20.32 9.65
Arm D 0.211 0.284 28.98 12.20
Arm E 0.899 0.033 47.72 10.62
Arm F 0.901 0.028 48.03 10.44
At least 1 0.982 0.723 - -
All 0.006 0.004 215.51 8.27
Table 3: (1) Probability of concluding a treatment arm, at least one treatment arm, or all treatment arms (arms B to F) to be effective at the end of the trial, (2) probability of stopping early a treatment arm, at least one treatment arm, or all treatment arms for futility (arms B to F), (3)(4) expected sample size per arm (A to F) and in total, and corresponding standard deviation, under the null and alternative hypotheses (refer to Table 2) in Case study 2.1
Figure 5: Barplot of the probability of stopping (y-axis) at each look (x-axis) as observed across the 10,000 simulated trials for the “Alternative” scenario for Case Study 2.1. The vertical blue dashed line shows the estimate of the expected sample size (ESS). As the trial only stops if all arms are dropped for futility, the expected sample size almost equals 216, the maximum sample size.
Figure 6: Violin plot (and boxplot) of the simulated sample sizes per group and corresponding expected sample sizes for the “Alternative” scenario for Case Study 2.1 across the 10,000 simulated trials. The sample size per group of each trial are displayed. For target parameters, these points are colour-coded and shaped by the decision and timing when the trial stopped. We can note that the ineffective arms ‘B’ and ‘C’ are are often dropped early for futility and that more patients are allocated to effective arms ‘E’ and ‘F’.
Figure 7: Scatterplot of target parameter log odds ratio estimates (posterior mean) versus sample size for each intervention arm (separate plots) at the time of stopping the arm, along with the decision at the time of stopping for the “Alternative” scenario for Case Study 2.1 across the 10,000 simulated trials. The timing of the stop is also provided (“Early stopping” = stopping at an interim analysis). We can note that log odds ratio estimates are typically negative when dropping arms early for futility.
4.1.2 Case Study 2.2
[…]
Conclusion Sample size
Group Efficacy(1) Futility(2) Average(3) Stand. Dev.(4)
Under the global null hypothesis:
Control (Arm A) - - 39.75 14.79
Arm B 0.030 0.965 22.98 13.52
Arm C 0.031 0.964 22.71 13.56
Arm D 0.030 0.965 22.46 13.49
Arm E 0.030 0.966 22.87 13.59
Arm F 0.029 0.965 22.94 13.78
At least 1 0.097 0.999 - -
All 0.000 0.883 153.71 45.82
Under the alternative hypothesis:
Control (Arm A) - - 53.89 10.09
Arm B 0.041 0.956 18.66 8.65
Arm C 0.037 0.960 18.47 8.54
Arm D 0.189 0.799 26.27 12.02
Arm E 0.863 0.126 48.32 12.13
Arm F 0.874 0.116 48.86 11.87
At least 1 0.968 0.993 - -
All 0.006 0.027 214.46 12.20
Table 4: (1) Probability of concluding a treatment arm, at least one treatment arm, or all treatment arms (arms B to F) are effective at the end of the trial, (2) probability of stopping early a treatment arm, at least one treatment arm, or all treatment arms for futility (arms B to F), (3)(4) expected sample size per arm (A to F) and in total, and corresponding standard deviation, under the null and alternative hypotheses (refer to Table 2) in Case study 2.2
Summary plots of the output for Case Study 2.2 can be found in Appendix C. The run time for these examples for a cluster, and standard Windows desktop are provided in the Appendix.
5 Discussion
Here we have presented an R package, BATSS (Couturier et al., 2025), which provides a modular, flexible, structured framework to simulate Bayesian adaptive trials. BATSS is an open source, script-based package that is extendable, and rich with options for adaptations for simulating Bayesian adaptive clinical trials. It allows for easy calculation of various trial operating characteristics and visualisation of some key metrics. We have provided an overview of the workflow and arguments required for the batss.glm function which can be used to simulate Bayesian adaptive trials whose outcome may be modelled by a GLM.
The BATSS package will help fill the gap between methodology research and implementation of Bayesian adaptive designs in practice. The BATSS package was designed to be accessible to R users with varying levels of programming expertise, and it is hoped that this package will reduce time spent on programming by avoiding the need to write simulation code from scratch for a range of Bayesian adaptive designs.
We believe that BATSS offers a good balance of flexibility by allowing the user to specify their RAR functions and algorithms for treatment allocations, as well as their early stopping/arm dropping rules, whilst taking care of the simulation work. This is an advantage over existing packages which may not offer sufficient flexibility in the scenarios and design settings that can be explored. A challenge of developing trial simulation software is to make it generalisable enough without overwhelming new users with the required inputs. Thus, there is a trade-off of flexibility and ease-of-use; to assist with this we provide several worked examples and functions for some commonly-used trial designs. Additionally, there are often implicit assumptions in the software code about the conduct of the clinical trial; we have tried to limit these assumptions by allowing user-specified functions and have described the assumptions that the batss.glm function makes.
Extending the BATSS package is fairly easy. This version of BATSS allows for the simulation of Bayesian adaptive designs which may be modelled using GLMs. We anticipate that additional functions will be added to the BATSS R package over time, with different functions for different types of analyses and designs. Future work will seek to extend it so that Bayesian adaptive designs may be simulated for time-to-event outcomes.
5.1 Limitations and Future Research
One of the main limitations with using R-INLA to perform the Bayesian approximations is that it is frequently being updated, which could potentially cause stability issues with the BATSS package (e.g., issues with implementing on cluster networks). Also, not all models are currently available in the INLA framework (e.g., heteroscedastic models) and so use of the BATSS package will be limited to the classes of models that are supported in the INLA framework (which appear to be sufficient for many commonly-used clinical trial applications).
The BATSS package does not currently support platform trial designs where treatment arms may be added or those implementing multi-factorial designs (with multiple domains of treatment). This is an area for future research. We also plan to extend the package to
  • support hierarchical models so that designs that implement information borrowing across subgroups can be implemented (e.g., basket trials). Hosseini et al. (2024) provide evidence that this is feasible and fast using INLA, and may provide more computational savings compared to MCMC for these models.
  • consider the enrollment times of patients or account for follow-up duration (leading to delays in data availability) or loss to follow-up and drop-out. Instead of assuming that analyses are conducted based on the number of participants completing follow-up (which, here, is also the same as the number recruited), a related extension would be to take into account the fact that some of the participants recruited have incomplete follow-up at the time of an interim analysis. Such features, typically not included in existing software for simulating Bayesian adaptive designs, would allow estimating further operating characteristics, mentioned by the guidelines described in Section 1.1, like, for example, the expected, minimum and maximum study times and number of recruited participants ($F^{19}_{2}$, $I^{20}_{1}$). We are currently extending BATSS to explicitly simulate accrual, drop-outs and account for the follow-up duration length (and therefore have participants with incomplete follow-up in the trial at interim analyses), and the next release of BATSS will contain these features. This is necessary in particular for survival/time-to-event endpoints.
  • propose point and interval estimators for the target parameters. Bias and mean squared error of point estimators as well as coverage of interval estimators are typically of interest for regulators ($F^{11}_{2}$, $I^{11}_{1}$, $I^{21}_{1}$. Whilst raw trial data for each simulated data set are available to users (by specifying extended = 2 in batss.glm) to define the properties of their planned estimators (see, e.g., Robertson et al. (2023a), Robertson et al. (2023b), Robertson et al. (2025a), Robertson et al. (2025b)), allowing this task to be performed with the same modular approach as to define other aspects of the trial, would be useful.
This version of BATSS is appropriate for slow recruiting trials and/or those that have endpoints that are observed over a short period (and no loss-to-follow-up).
Other extensions include extending the BATSS package for proportional odds models; trial designs where there is no control or reference group or where all treatment arms will be compared to one another (to determine the “best” treatment).